Find lim as h→0 f(x+h)-f(x)/h

a. f(x) = 1-x/x+2

b. f(x) = sqrt x²-3


1
Expert's answer
2021-11-03T10:06:26-0400

(a) "\lim\nolimits_h\to0\lbrack\frac{f(x+h)-f(x)}{h}\rbrack"


"=\lim\nolimits_h\to0\lbrack\frac{{\frac{1-(x+h)}{x+h+2}}-{\frac{1-x}{x+2}}}{h}\rbrack"


"=\lbrack\frac{{\frac{1-(x+h)}{x+h+2}}-{\frac{1-x}{x+2}}}{h}\rbrack"


"=\lim\nolimits_h\to0\lbrack\frac{((x+2)(1-x-h))-((1-x)(x+h+2))}{h(x+h+2)(x+2)}\rbrack"


"=\frac{3}{2\sqrt{-x+1}\left((x+2)^{\smash{\frac{3}{2}}}\right)}"



(b) the derivative of f(x) with respect to x is defined as:

"\lim\nolimits_h\to0\lbrack\frac{f(x+h)-f(x)}{h}\rbrack"


= "\lim\nolimits_h\to0\lbrack\frac{\sqrt{\left((x+h)^{\smash{2}}\right)-3)}-\sqrt{\left((x)^{\smash{2}}-3\right)}}{h}\rbrack"


"=\lbrack\frac{\sqrt{\left((x+h)^{\smash{2}}\right)-3)}-\sqrt{\left((x)^{\smash{2}}-3\right)}}{h}\rbrack"

"=\lim\nolimits_h\to0\lbrack\frac{\sqrt{\left((x+h)^{\smash{2}}\right)-3)}-\sqrt{\left((x)^{\smash{2}}-3\right)}}{h}\rbrack\lbrack\frac{\sqrt{\left((x+h)^{\smash{2}}\right)-3)}+\sqrt{\left((x)^{\smash{2}}-3\right)}}{\sqrt{\left((x+h)^{\smash{2}}\right)-3)+\sqrt{\left((x^{\smash{2}}\right)-3)}}}\rbrack"


"=\frac{x}{\sqrt{\left(x^{\smash{2}}\right)-3}}"